I am solving quadratic equations by completing the square:
9x^2-9x=18
I have tried dividing everything by 9 first, but I didn't know if I needed for that to be first or if I find half of the coefficient and square it first.

Answer by Alan3354(69443)

9x^2-9x=18
Divide by 9 first
x^2-x = 2
x^2-x + 1/4 = 9/4
(x - 1/2)^2 = (3/2)^2
etc